OpenStudy (anonymous):

2log6(x)=log6(2)+log6(3x+20) solve for x

OpenStudy (anonymous):

2log6(x) = log6(x)² Now, since all the terms have the same base 6, do you know what to do next?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

remove log

OpenStudy (anonymous):

the two log laws you need here are as follows: \[alog(b) = \log(b^a) \text{ , and } \log(a) + \log(b) = \log(ab)\]

OpenStudy (anonymous):

Great, could you show the step?

OpenStudy (anonymous):

x^2=2+3x+22 and then just solve for x?

OpenStudy (anonymous):

not quite, before you get rid of logs you need only one log on each side. like this: log(X) = log(Y) implies X =Y you need to group together the logs on the right using the second rule i posted :)

OpenStudy (anonymous):

loga + log b= log(ab)

OpenStudy (anonymous):

@nasryn I should have mentioned, please apply the log rule before get rid of log6

OpenStudy (anonymous):

a = 2, b = 3x + 20